Find the sum of all 15 terms in a sequence 7, 3, -1, -5 ...tn,
Found 2 solutions by luke94, richwmiller:
Answer by luke94(89)

firstly note your sequence is arithmetic as it is decreasing by -4.
write out your sequence: 7,3,-1,-5,-9,-13,-17,-21,-25,-29,-33,-37,-41,-45,-49
add them all up and your answer should be = -315

Answer by richwmiller(17219)

S = ½(a + an)n
S = ½(2a + (n-1)d)n
an = a + (n - 1)d

We will use
S = ½(2a + (n-1)d)n
since we don't know the 15th term
S = ½(2*7 + (14)(-4))15
S = (7 + (7)(-4))15
S = (7 -28)15
S = (-21)15
S=-315